# High-dimensional analysis of ridge regression for non-identically distributed data with a variance profile

### v2

@@ A.3 Proof of the main results
Since Assumption 2.2 hold true for the variance profile $\widetilde{S}_{p}$, it follows that $(\widetilde{S}_{p})_{p\geq 1}$ satisfy (A.3). Thus, we deduce from Theorem A.1 that
$\frac{1}{p}\operatorname{Tr}[Q_{p}(z)\widetilde{S}_{p}]\sim\frac{1}{p}\operatorname{Tr}[T_{p}(z)\widetilde{S}_{p}]\quad\mathrm{and}\quad\frac{1}{p}\operatorname{Tr}[Q_{p}^{\prime}(z)\widetilde{S}_{p}]\sim\frac{1}{p}\operatorname{Tr}[T^{\prime}(z)\widetilde{S}_{p}].$ (A.7)
@@
We assume that Assumptions 2.1 and 2.2 hold true in all this proof. According to [HLN07][Theorem 2.4], $T_{p}(-\lambda)$ can be expressed as follows for $\lambda>0$
$T_{p}(-\lambda)=\int_{\mathbb{R}^{+}}\frac{\mu(dw)}{w+\lambda},$ (A.8)
where $\mu=(\mu_{ij})$ is a positive $p\times p$ matrix valued measure such that $\mu_{ij}(\mathbb{R}^{+})=\delta_{ij}$. Moreover, $\lambda\mapsto\frac{1}{w/\lambda+1}$ is differentiable for $w\in\mathbb{R}^{+}$, its derivative $w\mapsto\frac{w}{(w+\lambda)^{2}}$ is measurable for $\lambda>0$ and $|\frac{w}{(w+\lambda)^{2}}|\leq 1$, thus $T_{p}(-\lambda)$ is differentiable and
$T_{p}^{\prime}(-\lambda)=\int_{\mathbb{R}^{+}}\frac{\mu(dw)}{(w+\lambda)^{2}}.$
As stated in Proposition 2.1 $\frac{1}{p}\mathrm{Tr}[T_{p}(z)]$ is a Stieltjes transform
$\frac{1}{p}\mathrm{Tr}[T_{p}(z)]=\int_{\mathbb{R}^{+}}\frac{\nu_{p}(dw)}{w+\lambda},$ (A.9)
where there exists $\pi_{\ast}\in[0,1]$ and $\pi:(0,+\infty)\rightarrow[0,+\infty)$ a locally Hölder-continuous function such that $\nu_{p}(dw)=\pi_{\ast}\delta_{0}(dw)+\pi(dw)\mathbf{1}_{w>0}dw$ [AEK17b][Theorem 2.1].
Under the additional Assumptions 2.3 and 2.4, [AEK17b][Theorem 2.9] asserts that there exists $\tau>0$ such that $\pi((0,\tau])=0$ and if $p<n$ then $\pi_{\ast}=0$. Hence we get from Equation (A.8) that
$\int_{0}^{\tau}\frac{\frac{1}{p}\mathrm{Tr}(\mu(dw))}{w+\lambda}=\int_{0}^{\tau}\frac{\nu_{p}(dw)}{w+\lambda}=0.$ (A.10)
Since $\frac{1}{w+\lambda}>0$ for $\lambda>0$ and $w\geq 0$, we deduce from Equation (A.10) that $\frac{1}{p}\sum_{i=1}^{p}\mu_{ii}((0,\tau])=0$. Which gives us $\mu_{ii}((0,\tau])=0$ for $1\leq i\leq p$ because $(\mu_{ii})_{1\leq i\leq p}$ are positive measures. Hence
$T_{p}(-\lambda)=\frac{\pi_{\ast}}{\lambda}+\int_{\tau}^{+\infty}\frac{\mu(dw)}{w+\lambda}.$
Let’s focus on the case $p<n$, we then have
$T_{p}(-\lambda)=\int_{\tau}^{+\infty}\frac{\mu(dw)}{w+\lambda}.$
@@
As stated in Section 2.3, in the case of a quasi doubly stochastic variance profile, $T_{p}(z)$ is a scalar matrix. Especially, $T_{p}(z)=m_{p}(z)I_{p}$ where $m_{p}(z)$ is the Stieltjes transform of the Marchenko-Pastur distribution. Moreover, following Equation (A.9) we get
$m_{p}(z)=\int_{\mathbb{R}^{+}}\frac{\nu_{p}(dw)}{w+\lambda},$
with $\nu_{p}(dw)=\pi_{\ast}\delta_{0}(dw)+\pi(dw)\mathbf{1}_{w>0}dw$. We have seen in the previous proof that there exists $\tau>0$ such that $\nu_{p}((0,\tau])=0$ and if $p<n$ then $\pi_{\ast}=0$. This implies that 00 belongs to the domain of definition of $m_{p}(.)$. Hence $\lim\limits_{\lambda\rightarrow 0}T_{p}(-\lambda)=\lim\limits_{\lambda\rightarrow 0}m_{p}(-\lambda)=m_{p}(0)$ and consequently $T_{p}(0^{-})=m_{p}(0)$.
@@
We assume that Assumptions 2.1 and 2.2 hold true in all this proof. We first focus on the limit with $\lambda\rightarrow+\infty$. According to [HLN07][Theorem 2.4], $T_{p}(-\lambda)$ can be expressed as follows for $\lambda>0$
@@
where $\mu=(\mu_{ij})$ is a $p\times p$ matrix valued measure such that $\mu(\mathbb{R}^{+})=I_{p}$. Since $w\mapsto\frac{1}{w/\lambda+1}$ is a measurable function for $\lambda>0$,$\lim\limits_{\lambda\rightarrow+\infty}\frac{1}{w/\lambda+1}=1$ for $w\in\mathbb{R}^{+}$ and $|\frac{1}{w/\lambda+1}|\leq 1$, we deduce from the dominated convergence theorem that
$\int_{\mathbb{R}^{+}}\frac{\mu(dw)}{w/\lambda+1}\xrightarrow[\lambda\rightarrow+\infty]{}I_{p}.$
Which finally gives us
@@
Moreover, $\lambda\mapsto\frac{1}{w/\lambda+1}$ is differentiable for $w\in\mathbb{R}^{+}$, its derivative $w\mapsto\frac{w}{(w+\lambda)^{2}}$ is measurable for $\lambda>0$ and $|\frac{w}{(w+\lambda)^{2}}|\leq 1$, thus $T_{p}(-\lambda)$ is differentiable and
$T_{p}^{\prime}(-\lambda)=\int_{\mathbb{R}^{+}}\frac{\mu(dw)}{(w+\lambda)^{2}}.$
We can prove the following limits the same way we proved (A.11)
@@
Hence by combining (A.11) and (A.12), we get the limit of the predictive risk for large $\lambda$:
$\lim\limits_{\lambda\rightarrow+\infty}r_{\lambda}(\Upsilon_{n})=\frac{\alpha^{2}}{p}\mathrm{Tr}[\widetilde{S}_{p}]+\sigma^{2}.$
Let’s now compute the limit of the predictive risk for small $\lambda$. Considering the definition of $T_{p}(0^{-})$ and $T_{p}^{\prime}(0^{-})$ given in Lemma 3.2, we directly get the limit of $r_{\lambda}(\Upsilon_{n})$ when $p<n$
$\lim_{\lambda\rightarrow 0,\;\lambda>0}r_{\lambda}(\Upsilon_{n})=\sigma^{2}+\frac{\sigma^{2}}{n}\mathrm{Tr}[\widetilde{S}_{p}T_{p}(0^{-})].$
Let’s now suppose that $p>n$. We have from (2.1) that
@@
$r_{\lambda}(\Upsilon_{n})=\sigma^{2}+\frac{\alpha^{2}}{p}\operatorname{Tr}[\widetilde{S}_{p}\kappa(\lambda)(\Sigma_{n}+\kappa(\lambda))^{-1}]+\left(\frac{\sigma^{2}}{n}-\frac{\lambda\alpha^{2}}{p}\right)Tr[\kappa^{\prime}(\lambda)\widetilde{S}_{p}\Sigma_{n}(\Sigma_{n}+\kappa(\lambda))^{-2}].$
@@
$\lim_{\lambda\rightarrow 0,\;\lambda>0}r_{\lambda}(\Upsilon_{n})=\frac{\alpha^{2}}{p}\mathrm{Tr}[\widetilde{S}_{p}\kappa(0^{+})(\Sigma_{n}+\kappa(0^{+}))^{-1}]+\frac{\sigma^{2}}{n}\mathrm{Tr}[\kappa^{\prime}(0^{+})\widetilde{S}_{p}\Sigma_{n}(\Sigma_{n}+\kappa(0^{+}))^{-2}]+\sigma^{2}.$
@@
Let us prove Equation (3.5). The function $g:\lambda\mapsto\hat{r}_{\lambda}(X_{n})$ is differentiable for $\lambda>0$, with
$g^{\prime}(\lambda)=2\Big{(}\frac{\lambda\alpha^{2}}{p}-\frac{\sigma^{2}}{n}\Big{)}\mathrm{Tr}[\Sigma_{n}\widehat{\Sigma}_{n}Q_{p}(\lambda)^{3}]$
Since the symmetric and positive semi-definite matrices $\widehat{\Sigma}_{n}$ and $Q_{p}(\lambda)^{3}$ are commuting, then $\Sigma_{n}\hat{\Sigma}_{n}Q_{p}(\lambda)^{3}$ is also symmetric positive semi-definite.
Thus $\mathrm{Tr}[\Sigma_{n}\widehat{\Sigma}_{n}Q_{p}(\lambda)^{3}]\geq 0$ which means that the sign of $g^{\prime}(\lambda)$ only depends on $\Big{(}\frac{\lambda\alpha^{2}}{p}-\frac{\sigma^{2}}{n}\Big{)}$. Indeed, if $\lambda\leq\lambda_{\ast}$ then $g^{\prime}(\lambda)\leq 0$, and if $\lambda\geq\lambda_{\ast}$ then $g^{\prime}(\lambda)\geq 0$. This proves that Equation (3.5) holds true, and this concludes the proof of Corollary 3.3. ∎
@@ A.4 Random matrices with a variance profile in free probability
$\mathcal{R}_{\Sigma_{n}}^{\square}(\mathcal{Z})=\mathrm{deg}\left(\frac{1}{n}\Gamma_{n}^{t}\left[I_{n}-\mathrm{deg}\left(\frac{1}{n}\Gamma_{n}\mathcal{Z}\right)\right]^{-1}\right),\quad\mbox{for all}\quad\mathcal{Z}\in\mathbb{D}_{p}(\mathbb{C})^{-}.$

### v3

@@ A.3 Proof of the main results
Since Assumption 1.3 hold true for the variance profile $\widetilde{S}_{p}$ and $\Upsilon_{n}$, it follows that $(\widetilde{S}_{p})_{p\geq 1}$ and $(\widehat{\Sigma}_{n})_{n\geq 1}$ satisfy (3.3). Moreover the family of identity matrices, namely $(I_{p})_{p\geq 1}$, also satisfies (3.3) thus, we deduce from Theorem 3.1 that
$\frac{1}{p}\operatorname{Tr}[Q_{p}(z)\widetilde{S}_{p}]\sim\frac{1}{p}\operatorname{Tr}[T_{p}(z)\widetilde{S}_{p}]$ , $\frac{1}{p}\operatorname{Tr}[Q_{p}^{\prime}(z)\widetilde{S}_{p}]\sim\frac{1}{p}\operatorname{Tr}[T^{\prime}(z)\widetilde{S}_{p}]$ (A.3)
$\frac{1}{p}\operatorname{Tr}[Q_{p}^{\prime}(z)\widehat{\Sigma}_{n}]\sim\frac{1}{p}\operatorname{Tr}[T_{p}^{\prime}(z)\widehat{\Sigma}_{n}]$ , $\frac{1}{p}\operatorname{Tr}[Q_{p}^{\prime}(z)]\sim\frac{1}{p}\operatorname{Tr}[T^{\prime}(z)].$ (A.4)
@@
We assume that Assumptions 1.2 and 1.3 hold true in all this proof. According to [HLN07][Theorem 2.4], one has that $T_{p}(-\lambda)=\int_{\mathbb{R}^{+}}\frac{\mu(dw)}{w+\lambda}$ for $\lambda>0$, where $\mu=(\mu_{ij})$ is a positive $p\times p$ matrix valued measure such that $\mu_{ij}(\mathbb{R}^{+})=\delta_{ij}$. Moreover, $\lambda\mapsto\frac{1}{w/\lambda+1}$ is differentiable for $w\in\mathbb{R}^{+}$, its derivative $w\mapsto\frac{w}{(w+\lambda)^{2}}$ is measurable for $\lambda>0$ and $|\frac{w}{(w+\lambda)^{2}}|\leq 1$, thus $T_{p}(-\lambda)$ is differentiable and $T_{p}^{\prime}(-\lambda)=\int_{\mathbb{R}^{+}}\frac{\mu(dw)}{(w+\lambda)^{2}}$.
As stated in Proposition 2.1 $\frac{1}{p}\mathrm{Tr}[T_{p}(z)]$ is a Stieltjes transform $\frac{1}{p}\mathrm{Tr}[T_{p}(z)]=\int_{\mathbb{R}^{+}}\frac{\nu_{p}(dw)}{w+\lambda}$, where there exists $\pi_{\ast}\in[0,1]$ and $\pi:(0,+\infty)\rightarrow[0,+\infty)$ a locally Hölder-continuous function such that $\nu_{p}(dw)=\pi_{\ast}\delta_{0}(dw)+\pi(dw)\mathbf{1}_{w>0}dw$ [AEK17b][Theorem 2.1].
Under the additional Assumptions 1.4 and 1.5, [AEK17b][Theorem 2.9] asserts that there exists $\tau>0$ such that $\pi((0,\tau])=0$ and if $p<n$ then $\pi_{\ast}=0$. Since $T_{p}(-\lambda)=\int_{\mathbb{R}^{+}}\frac{\mu(dw)}{w+\lambda}$ for $\lambda>0$, one get that $\int_{0}^{\tau}\frac{\frac{1}{p}\mathrm{Tr}(\mu(dw))}{w+\lambda}=\int_{0}^{\tau}\frac{\nu_{p}(dw)}{w+\lambda}=0$.
Since $\frac{1}{w+\lambda}>0$ for $\lambda>0$ and $w\geq 0$, we deduce from the previous equation that $\frac{1}{p}\sum\limits_{i=1}^{p}\mu_{ii}((0,\tau])=0$. Which gives us $\mu_{ii}((0,\tau])=0$ for $1\leq i\leq p$ because $(\mu_{ii})_{1\leq i\leq p}$ are positive measures. Hence $T_{p}(-\lambda)=\frac{\pi_{\ast}}{\lambda}+\int_{\tau}^{+\infty}\frac{\mu(dw)}{w+\lambda}$.
Let’s focus on the case $p<n$, we then have $T_{p}(-\lambda)=\int_{\tau}^{+\infty}\frac{\mu(dw)}{w+\lambda}$.
@@
As stated in Section 2.3, in the case of a quasi doubly stochastic variance profile, $T_{p}(z)$ is a scalar matrix. Especially, $T_{p}(z)=m_{p}(z)I_{p}$ where $m_{p}(z)$ is the Stieltjes transform of the Marchenko-Pastur distribution. Moreover since $\frac{1}{p}\mathrm{Tr}[T_{p}(z)]=\int_{\mathbb{R}^{+}}\frac{\nu_{p}(dw)}{w+\lambda}$, we get $m_{p}(z)=\int_{\mathbb{R}^{+}}\frac{\nu_{p}(dw)}{w+\lambda},$ with $\nu_{p}(dw)=\pi_{\ast}\delta_{0}(dw)+\pi(dw)\mathbf{1}_{w>0}dw$. We have seen in the previous proof that there exists $\tau>0$ such that $\nu_{p}((0,\tau])=0$ and if $p<n$ then $\pi_{\ast}=0$. This implies that 00 belongs to the domain of definition of $m_{p}(.)$. Hence $\lim\limits_{\lambda\rightarrow 0}T_{p}(-\lambda)=\lim\limits_{\lambda\rightarrow 0}m_{p}(-\lambda)=m_{p}(0)$ and consequently $T_{p}(0^{-})=m_{p}(0)$.
@@
We assume that Assumptions 1.2 and 1.3 hold true in all this proof. We first focus on the limit with $\lambda\rightarrow+\infty$. According to [HLN07][Theorem 2.4], $T_{p}(-\lambda)$ can be expressed as follows for $\lambda>0$
@@
where $\mu=(\mu_{ij})$ is a $p\times p$ matrix valued measure such that $\mu(\mathbb{R}^{+})=I_{p}$. Since $w\mapsto\frac{1}{w/\lambda+1}$ is a measurable function for $\lambda>0$,$\lim\limits_{\lambda\rightarrow+\infty}\frac{1}{w/\lambda+1}=1$ for $w\in\mathbb{R}^{+}$ and $|\frac{1}{w/\lambda+1}|\leq 1$, we deduce from the dominated convergence theorem that $\int_{\mathbb{R}^{+}}\frac{\mu(dw)}{w/\lambda+1}\xrightarrow[\lambda\rightarrow+\infty]{}I_{p}.$ Which finally gives us
@@
Moreover, $\lambda\mapsto\frac{1}{w/\lambda+1}$ is differentiable for $w\in\mathbb{R}^{+}$, its derivative $w\mapsto\frac{w}{(w+\lambda)^{2}}$ is measurable for $\lambda>0$ and $|\frac{w}{(w+\lambda)^{2}}|\leq 1$, thus $T_{p}(-\lambda)$ is differentiable and $T_{p}^{\prime}(-\lambda)=\int_{\mathbb{R}^{+}}\frac{\mu(dw)}{(w+\lambda)^{2}}.$ We can prove the following limits the same way we proved (A.5)
@@
Hence by combining (A.5) and (A.6), we get the limit of the predictive risk for large $\lambda$, $\lim\limits_{\lambda\rightarrow+\infty}r_{\lambda}^{test}(X_{n})=\frac{\alpha^{2}}{p}\mathrm{Tr}[\widetilde{S}_{p}]+\sigma^{2}.$ Let’s now compute the limit of the predictive risk for small $\lambda$. Considering the definition of $T_{p}(0^{-})$ and $T_{p}^{\prime}(0^{-})$ given in Lemma 3.2, we directly get the limit of $r_{\lambda}^{test}(X_{n})$ when $p<n$, $\lim_{\lambda\rightarrow 0,\;\lambda>0}r_{\lambda}^{test}(X_{n})=\sigma^{2}+\frac{\sigma^{2}}{n}\mathrm{Tr}[\widetilde{S}_{p}T_{p}(0^{-})].$ Let’s now suppose that $p>n$. We have from (2.1) that
@@
$r_{\lambda}^{test}(X_{n})$ $=$ $\sigma^{2}+\frac{\alpha^{2}}{p}\operatorname{Tr}[\widetilde{S}_{p}\kappa(\lambda)(\Sigma_{n}+\kappa(\lambda))^{-1}]$
$+\left(\frac{\sigma^{2}}{n}-\frac{\lambda\alpha^{2}}{p}\right)Tr[\kappa^{\prime}(\lambda)\widetilde{S}_{p}\Sigma_{n}(\Sigma_{n}+\kappa(\lambda))^{-2}].$
@@
$\lim_{\lambda\rightarrow 0,\;\lambda>0}r_{\lambda}^{test}(X_{n})$ $=$ $\frac{\alpha^{2}}{p}\mathrm{Tr}[\widetilde{S}_{p}\kappa(0^{+})(\Sigma_{n}+\kappa(0^{+}))^{-1}]$
$+\frac{\sigma^{2}}{n}\mathrm{Tr}[\kappa^{\prime}(0^{+})\widetilde{S}_{p}\Sigma_{n}(\Sigma_{n}+\kappa(0^{+}))^{-2}]+\sigma^{2}.$
@@
Let us prove Corollary 3.3. The function $g(\lambda)=\hat{r}_{\lambda}^{test}(X_{n})$ is differentiable for $\lambda>0$, with $g^{\prime}(\lambda)=2\Big{(}\frac{\lambda\alpha^{2}}{p}-\frac{\sigma^{2}}{n}\Big{)}\mathrm{Tr}[\Sigma_{n}\widehat{\Sigma}_{n}Q_{p}(\lambda)^{3}].$ Since the symmetric and positive semi-definite matrices $\widehat{\Sigma}_{n}$ and $Q_{p}(\lambda)^{3}$ are commuting, then $\Sigma_{n}\hat{\Sigma}_{n}Q_{p}(\lambda)^{3}$ is also symmetric positive semi-definite.
Thus $\mathrm{Tr}[\Sigma_{n}\widehat{\Sigma}_{n}Q_{p}(\lambda)^{3}]\geq 0$ which means that the sign of $g^{\prime}(\lambda)$ only depends on $\Big{(}\frac{\lambda\alpha^{2}}{p}-\frac{\sigma^{2}}{n}\Big{)}$. Indeed, if $\lambda\leq\lambda_{\ast}$ then $g^{\prime}(\lambda)\leq 0$, and if $\lambda\geq\lambda_{\ast}$ then $g^{\prime}(\lambda)\geq 0$. This proves that Corollary 3.3 holds true, and this concludes the proof of Corollary 3.3. ∎
@@ A.4 Random matrices with a variance profile in free probability
$\mathcal{R}_{\Sigma_{n}}^{\square}(\mathcal{Z})=\mathrm{deg}\left(\frac{1}{n}\Gamma_{n}^{\top}\left[I_{n}-\mathrm{deg}\left(\frac{1}{n}\Gamma_{n}\mathcal{Z}\right)\right]^{-1}\right),\quad\mbox{for all}\quad\mathcal{Z}\in\mathbb{D}_{p}(\mathbb{C})^{-}.$